# Poisoning with A Pill: Circumventing Detection in Federated Learning

### v1

@@ 11       $\mathop{param}\leftarrow\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\}$;
12       $\Delta\boldsymbol{\widetilde{g}}_{t+1}\leftarrow$ Estimation($i$,$t$,$\boldsymbol{g}_{t}$,$\mathop{param}$);
@@ 16       $\mathop{param}=\mathop{param}\bigcup\{\boldsymbol{M}_{all}=M+M_{disc}\}$;
17       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$SimAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
18       $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow$DistAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
@@
Algorithm 1 Our augmentation method’s workflow. Without losing generality, we assume the first $m$ clients as malicious clients.
5.1 Overview of Our Method
Fig. 1 presents the three key stages. Table I presents notations of main symbols in this paper.
[…]
We are the first that propose a universal attack augmentation pipeline for most FL poisoning attacks, considering stealthiness, compatibility, and generality. The detailed workflow of our method is shown in Algorithm 1.
5.2 Pill Construction
@@
This stage aims to construct a pill structure for augmenting the stealthiness while retaining the attack effectiveness before being augmented. The pill is carefully crafted to involve a minimal subset of parameters from specific positions of the target model. We first define a pill’s blueprint as the pill’s graphic structure, independent of target model parameters. Then, we propose a dynamic pill search algorithm to identify and map concrete parameters from the target model to the blueprint.
Designing Pill Blueprint. The blueprint design is inspired by SRA [15], which shows that poisoning a narrow subnetwork (one neuron/channel in each layer) is adequate to effectively inject backdoors into machine learning models (not in the FL setting). However, their technique cannot be used for our purposes as their subnet architecture is very specific. It does not support attacking various targets; it is a fixed and pre-selected subnet without considering the dynamics of model training in FL; and its poisoned subnet is not stealthy, having substantially larger weight values compared to others due to the need to disseminate the poison effect through such a small pre-selected network. Therefore, we propose a novel blueprint method, in which the subnet structure is general, and its instantiations (i.e., the concrete subnets) vary across steps in the FL training procedure, including important neurons by a dynamic search algorithm. This allows small weight changes because poisoning important neurons enables easy dissemination, maximizing attack stealthiness. In particular, the pill blueprint is designed to accommodate various target classes of different FL poisoning attacks. It achieves this by manipulating the outputs relevant to multiple classes simultaneously, via disrupting all the output neurons together. Hence, our pill blueprint design follows the rules below:
  1. 1.
    The pill blueprint only contains one neuron in each linear layer or one channel in each convolutional layer, except for the last two layers. Suppose $\mathcal{N}_{i}^{p}$ represents the neuron/channel number in Layer $i$ in our pill blueprint, then $\mathcal{N}_{i}^{p}=1$ when $i<L-1$, where $L$ is the total layer counts in our pill blueprint.
  2. 2.
Dynamic Pill Search. According to existing studies on neural network pruning [40, 41, 42, 43, 44], parameters with a larger magnitude typically dominate the model’s performance. The optimal solution is hence to examine the model parameters to search for a globally optimal pill that encompasses the most important parameters.
However, such a globally optimal pill could be identified via a pruning-based method [51, 52], and hence our attack could be easily detected. Besides, searching for a globally optimal pill is inefficient when the model has a large number of parameters. Thus, we search for an approximate pill instead, with an attacker-defined start point, and only evaluate a small subset of the entire model’s parameters. We name the search algorithm as “approximate max pill search”. A complete procedure of this search algorithm consists of the following four steps. To improve readability, we use “neuron” to represent both neurons in fully connected layers and channels in convolutional layers, and we use the classification task as an example:
@@
After visiting $l_{L-1}$ layers, where $L$ is the total number of layers in the target model, $||\mathcal{V}_{L-1}||$ should equal to the neuron number in $l_{L}$ of the target model, which also equals to the number of classes. We select all the neurons in the target model’s $l_{L}$ layer into our pill to construct $\mathcal{V}_{L}$. Then, we only record the parameters from one neuron in $\mathcal{V}_{L-1}$ to only one neuron in $\mathcal{V}_{L}$ based on the index order (i.e., the first neuron in $\mathcal{V}_{L-1}$ is paired with the first neuron in $\mathcal{V}_{L}$). Since $||\mathcal{V}_{L-1}||$ equals to $||\mathcal{V}_{L}||$, the number of recorded parameters equals to the number of classes, avoiding poisoning too many parameters in a single layer.
@@
With the recorded $\mathcal{V}_{i}$ and the corresponding parameters, we construct two masks $\boldsymbol{M}$ and $\boldsymbol{M}_{disc}$. The mask $\boldsymbol{M}$ records the pill’s parameters in the target model, and the disconnection mask $\boldsymbol{M}_{disc}$ records the parameters of the connections between the pill and the rest of the target model. $\boldsymbol{M}$ is used for poisoning, while $\boldsymbol{M}_{disc}$ is used to disconnect the poison pill from the target model, maintaining the integrity and performance of the pill during poisoning. The two masks have the same shape as the target model’s parameters. To construct $\boldsymbol{M}$, we set the locations corresponding to the pill parameters to $1$, and the others to 00. Based on $\boldsymbol{M}$, we can similarly obtain the corresponding disconnected mask $\boldsymbol{M}_{disc}$, which sets the locations corresponding to the parameters from neurons other than $\mathcal{V}_{i-1}$ to $\mathcal{V}_{i}$ in each model layer $l_{i}$ (except for the $L$th layer since we choose all the neurons from it), and also those corresponding to parameters from $\mathcal{V}_{i-1}$ to other pill irrelevant neurons in each model layer $l_{i}$ to $1$. The two masks are used in the Pill Injection Stage.
Example. Fig. 2 presents a concrete example of the search algorithm in a 4-layer linear model. Since the start point is randomly selected by the attacker, defense methods can hardly guess it without any prior knowledge. In the example, suppose the model is a 4-layer linear models for a binary classification task. Then the pill blue print contains one neuron in the first two layers, and contains two neurons in the last two layers. Initially, we randomly select a start neuron, specifically the second neuron in the first layer in the example. Then, we conduct layer-wise searching when visiting the second and third layer, selecting the parameters with the highest magnitudes. At the forth layer (output layer), we pair the two output neurons with the two selected neurons in the third layer based on the index order. And we finally construct two pill-related masks accordingly.
@@
TABLE II: Different dynamic patterns in our augmentation method, utilizing along with the max subnetwork searching.
@@
$\textsc{Pattern}_{1}$ All layers use adaptive searching strategy
@@
$\textsc{Pattern}_{2}$ All layers use one-time searching strategy
@@
$\textsc{Pattern}_{3}$Centerstack[l]FE uses with adaptive searching strategy
CLS uses repeated searching strategy
$\textsc{Pattern}_{4}$Centerstack[l]FE uses repeated searching strategy
CLS uses with adaptive searching strategy
@@
$\textsc{Pattern}_{5}$Centerstack[l]FE uses with adaptive searching strategy
CLS uses with one-time searching strategy
$\textsc{Pattern}_{6}$Centerstack[l]FE uses with one-time searching strategy
CLS uses with adaptive searching strategy
@@
To make the pill search process dynamic, we also design several patterns to adaptively determine whether to change the pill in the training period, shown in Table II (FE represents the convolutional layers, CLS represents the linear layers). For more details about each specific dynamic pattern, please refer to Appendix C. The combination of the “approximate max pill search” with different dynamic patterns construct the complete dynamic pill search, considering both the stealthiness and the efficiency.
Figure 3: Cosine similarities between FLTrust server’s model update and malicious model update when malicious clients use different extra training rounds.
5.3 Pill Poisoning
The motivation to use an extra-trained model update as the reference model update is shown in Fig. 3. As shown in the figure, with the increasing number of extra training rounds on the malicious clients, the generated malicious model update becomes less opposite to the FLTrust [21] server’s model update. Thus, we adopt the extra training in our method and limit the extra training epoch number $E_{extra}$ to less than the number of malicious clients $m$ times the benign local training epoch number $E$, denoted as $E_{extra}\leq m\cdot E$. With this extra training epoch number limit, we do not violate the threat model since the attacker can utilize the data and computational resources of all the compromised clients.
5.4 Pill Injection
Pill Insertion & Disconnection. In this part, our goal is to insert the pill into the model, and minimize the impact of the benign model updates on our pill. We use an estimated global model update as the benign model update, which is estimated as the coordinate-wise mean values of all the normal model updates from the compromised clients. The estimation process in the Estimation() in Algorithm 1 Line 11 is hence presented as Equations (2),
where $\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}$ is the normal updates from the compromised clients. By aggregating information from multiple malicious clients, the estimated benign global model update is more similar to the genuine one, providing more budget for our poison pill.
After obtaining the estimated global model update $\boldsymbol{\Delta\widetilde{g}}_{t+1}$, we directly replace the parameters corresponding to the pill parameters (which have been poisoned in the previous stage) via the pill’s mask $\boldsymbol{M}$. Then, we replace the parameters that connect the pill and the other estimated global model updates with the disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$, using the disconnection mask $\boldsymbol{M}_{disc}$. The disconnection update $\Delta\boldsymbol{g}_{t+1}^{zero}$ is calculated as $0-\boldsymbol{g}_{t}$, and is bounded by the maximum and minimum values of the reference model update $\Delta\boldsymbol{\hat{g}^{m}_{t+1}}$. The disconnection update can gradually change the parameters of the connections between the pill and rest of the model to 00, and finally isolates the poison pill from the global model, guaranteeing the attacking effects of the poison pill.
1 function SimAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
@@ 5.4 Pill Injection
4       $\mathop{iter}\leftarrow 0$;
5       while Sim($\Delta\boldsymbol{\widetilde{g}}_{t+1},\Delta\boldsymbol{g}_{t+1}^{(i)}$)<$S_{max}$& $\mathop{iter}<C_{iter}$ do
@@ 12                   $\mathop{iter}\leftarrow\mathop{iter}+1$;
17
18function DistAdjust($\mathop{param}$,$\boldsymbol{\Delta\widetilde{g}}_{t+1}$,$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
@@
20       $\mathop{Dist}_{max}\leftarrow\mathop{max}\{||\Delta\boldsymbol{g^{\prime}}^{(i)}_{t+1}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||;i\in\{1,\cdots,m\}\}$;
21       $\mathop{Dist}\leftarrow||\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||$;
22       if $||C_{\downarrow}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||<||C_{\uparrow}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||$ then
@@ 26             $C_{dist}\leftarrow C_{\uparrow}$;
27
28      while $\mathop{Dist}\geq\mathop{Dist}_{max}$& $||C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||\leq\mathop{Dist}$ do
@@ 29             $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
Pill Adjustment. After the injection, we use a two-step adjustment to further adjust the pill, improving the generality against multiple detection metrics simultaneously. In this stage, we consider two prevailing detection metrics – distance and cosine similarity. To increase the cosine similarity between the poisoned model update and the benign model update in our method, we balance the magnitudes of both the poison pill’s parameters and the other benign parameters. Similarly, to minimize the distance discrepancy between the poisoned and benign model updates, we adjust the magnitude of the entire poisoned model update. Thus, we first use the similarity-based adjustment, then use the distance-based adjustment to balance the effectiveness and stealthiness poisoned model update. This two-step adjustment is particularly effective when combined with our method, which selectively poisons only a tiny subset of the model’s parameters. By altering just a few parameters, our method preserves a substantial number of benign parameters, which are crucial for making effective adjustments. As a result, the poisoned model update can bypass a wide range of defenses since they are typically designed based on the combination or variants of distance and cosine similarity metrics, and they usually do not anticipate such a focused and minimal interference in the model parameters. The details of the two seamless adjustments are as follows:
Similarity-based Adjustment.
@@
As shown in Line 1--12 of Algorithm 2, we first compute the maximum cosine similarity $S_{max}$ between the normal model updates from the compromised clients and the estimated global model update in the current round. Then, we iteratively and alternately reduce the magnitude of the poison pill’s parameters with the down-scaling factor $C_{\downarrow}$, and increase the magnitude of the rest estimated global model update’s parameters with the up-scaling factor $C_{\uparrow}$, until the cosine similarity between the entire poisoned model update and the estimated global model update is greater than $S_{max}$ or the adjustment total iteration is greater than the threshold $C_{iter}$.
Distance-based Adjustment.
@@
In the Distance-based Adjustment (Line 13--24 of Algorithm 2), we reuse the up-scaling factor $C_{\uparrow}$ and the down-scaling factor $C_{\downarrow}$ to adjust the magnitude of the entire poisoned model update. The intuition behind the Distance-based Adjustment is shown in Fig. 4. We first calculate the maximum distance between the normal model updates from the compromised clients and the estimated global model update in the current round. We use this maximum distance $Dist_{max}$ as the threshold in the distance-based adjustment. Then, we further determine the scaling factor that should be used by applying the two scaling factors $C_{\uparrow}$ and $C_{\downarrow}$ separately to the poisoned model update $\Delta\boldsymbol{g}_{t+1}^{(i)}$. The scaling factor that reduces the distance between $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is chosen as the initial scaling factor in the subsequent iterative scaling. We stop the scaling until the distance between the $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is smaller than $Dist_{max}$, or the scaling factor begins to increase such distance (reach the limit of the scaling).
6 Evaluation
This section assesses how our method enhances the effectiveness of existing FL poisoning attacks from multiple perspectives. We begin by evaluating the Augmentation Effectiveness of our method against four FL poisoning attacks, using nine prevailing defenses across three datasets, detailed in Section 6.2. Subsequently, we visualize the Stealthiness of our method under two prevailing detection metrics, as discussed in Section 6.3. Lastly, an Generality Analysis of our method is presented, which includes tests on various proportions of malicious clients, tests on both cross-silo and cross-device settings, and evaluates the impact of different pill search rules, outlined in Section 6.4. Our method significantly enhances the capabilities of existing FL poisoning attacks, successfully bypassing all $9$ baseline defenses in over 90% of cases, and increasing the error rates by up to seven times compared to the original attacks’ error rates. Moreover, it demonstrates robustness across varying data distributions, model architectures, proportions of malicious clients, and pill search rules.
6.1 Evaluation Settings
Attack, Defense, and Framework settings. In our experiments, we typically set the malicious proportion to $20\%$. We implement $9$ baseline aggregation rules, including FedAvg, FLTrust, Multi-Krum, Bulyan, Median, Trim, FLDetector, DnC, and Flame. We use our method to augment $4$ existing model poisoning attacks, including sign-flipping attack, Trim attack, Krum attack, and Min-Max attack. These attacks are chosen for their representativeness in illustrating the effectiveness of our method. We configure a 50-client FL system for both the MNIST and Fashion-MNIST datasets. For the CIFAR-10 dataset, a 30-client FL system is used. Our framework accommodates both cross-silo and cross-device settings. The entire framework is based on PyTorch [53].

### v2

@@ 11 $\mathop{param}\leftarrow\{\Delta\boldsymbol{g^{\prime}}^{(1),\cdots,(m)}_{t+1}\}$;
[…]
@@ 16 $\mathop{param}=\mathop{param}\bigcup\{\boldsymbol{M}_{all}=M+M_{disc}\}$;
17 $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrowSimAdjust$($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
18 $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrowDistAdjust$($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$);
@@
Algorithm 1 Our method’s workflow. Assume the first $m$ clients as malicious clients.
Appendix C Main Notations and Pseudo-Code
Table 6 presents the main notation used in the paper,W and Algorithm 1 shows the pseudo-code of our proposed attack augmentation method.
Appendix D Additional Details of Concrete Pill Blueprints
Table 4 and Table 5 illustrate the model structures of the CNN model and the simplified AlexNet, with their corresponding pill’s blueprints.
Appendix E Detailed Pill Search Algorithm
@@
A complete procedure of the “approximate max pill search” algorithm consists of the following four steps. To improve readability, we use “neuron” to represent both neurons in fully connected layers and channels in convolutional layers, and we use the classification task as an example:
@@
After visiting $l_{L-1}$ layers, where $L$ is the total number of layers in the target model, $\mathcal{V}_{L-1}||$ should equal to the neuron number in $l_{L}$ of the target model, which also equals to the number of classes. We select all the neurons in the target model’s $l_{L}$ layer into our pill to construct $\mathcal{V}_{L}$. Then, we only record the parameters from one neuron in $\mathcal{V}_{L-1}$ to only one neuron in $\mathcal{V}_{L}$ based on the index order (i.e., the first neuron in $\mathcal{V}_{L-1}$ is paired with the first neuron in $\mathcal{V}_{L}$). Since $\mathcal{V}_{L-1}||$ equals to $\mathcal{V}_{L}||$, the number of recorded parameters equals to the number of classes, avoiding poisoning too many parameters in a single layer.
@@
With the recorded $\mathcal{V}_{i}$ and the corresponding parameters, we construct two masks $\boldsymbol{M}$ and $\boldsymbol{M}_{disc}$. The mask $\boldsymbol{M}$ records the pill’s parameters in the target model, and the disconnection mask $\boldsymbol{M}_{disc}$ records the parameters of the connections between the pill and the rest of the target model. $\boldsymbol{M}$ is used for poisoning, while $\boldsymbol{M}_{disc}$ is used to disconnect the poison pill from the target model, maintaining the integrity and performance of the pill during poisoning. The two masks have the same shape as the target model’s parameters. To construct $\boldsymbol{M}$, we set the locations corresponding to the pill parameters to $1$, and the others to 0. Based on $\boldsymbol{M}$, we can similarly obtain the corresponding disconnected mask $\boldsymbol{M}_{disc}$, which sets the locations corresponding to the parameters from neurons other than $\mathcal{V}_{i-1}$ to $\mathcal{V}_{i}$ in each model layer $l_{i}$ (except for the $L$th layer since we choose all the neurons from it), and also those corresponding to parameters from $\mathcal{V}_{i-1}$ to other pill irrelevant neurons in each model layer $l_{i}$ to $1$. The two masks are used in the Pill Injection Stage.
Example. Figure 5 presents a concrete example of the search algorithm in a 4-layer linear model. Since the start point is randomly selected by the attacker, defense methods can hardly guess it without any prior knowledge. In the example, suppose the model is a 4-layer linear model for a binary classification task. Then the pill blue print contains one neuron in the first two layers, and contains two neurons in the last two layers. Initially, we randomly select a start neuron, specifically the second neuron in the first layer in the example. Then, we conduct layer-wise searching when visiting the second and third layers, selecting the parameters with the highest magnitudes. At the forth layer (output layer), we pair the two output neurons with the two selected neurons in the third layer based on the index order. And we finally construct two pill-related masks accordingly.
@@
$\textsc{Pattern}_{1}$ All layers use the adaptive searching strategy
$\textsc{Pattern}_{2}$ All layers use the one-time searching strategy
@@
$\textsc{Pattern}_{3}$ CLS layers use the repeated searching strategyFE layers use the adaptive searching strategy
@@
$\textsc{Pattern}_{4}$ CLS layers use the adaptive searching strategyFE layers use repeated searching strategy
@@
$\textsc{Pattern}_{5}$ CLS layers use the one-time searching strategyFE layers use the adaptive searching strategy
@@
$\textsc{Pattern}_{6}$ CLS layers use the adaptive searching strategyFE layers use the one-time searching strategy
Table 7: Different dynamic patterns in our method, utilizing along with “approximate max pill search”.
@@
To make the pill search process dynamic, we also design several patterns to adaptively determine whether to change the pill in the training period, shown in Table 7 (FE represents the convolutional layers, CLS represents the linear layers). For more details about each specific dynamic pattern, please refer to Appendix F. The combination of the “approximate max pill search” with different dynamic patterns constructs the complete dynamic pill search, considering both the stealthiness and the efficiency.
Appendix F Additional Details of The Dynamic Patterns in Our Method
Considering advantages and disadvantages of both the one-time searching strategy and the repeated searching strategy, we design a more flexible searching strategy, termed as “adaptive searching strategy”. In the adaptive searching strategy, our method searches the new pill only when the pill is not successfully injected into the global model in the last round. The condition:
$\texttt{Sim(}\boldsymbol{M}\odot\Delta\boldsymbol{g}_{t},\boldsymbol{M}\odot\Delta\boldsymbol{g}_{t}^{(i)}\texttt{)}<C_{search}$ (3)
should be satisfied to trigger the new subnetwork searching on malicious client $i$, where $C_{search}$ is set as $0.94$ in the experiments. The adaptive searching strategy is a more moderate version of repeated searching.
Since the three searching strategies have their unique advantages, we investigate different combinations of them in the experiments. We further divide the neural network into Feature Extractor (FE) and Classifier (CLS). Refer to the CNN model we used, the convolutional layers are regarded as FE, and the linear layers are regarded as CLS. We use different searching strategies in FE and CLS, respectively. In all the nine combinations, we test and keep six of them, noted as Pattern1 to Pattern6, shown in Table 7 in § 4.2. Such six patterns construct the entire dynamic pattern set used in our method.
Appendix G Additional Details of Pill Adjustment
1
function SimAdjust($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
@@ Appendix G Additional Details of Pill Adjustment
3 $S_{max}\leftarrow\mathop{max}(0,\mathop{max}\{$Sim($\Delta\boldsymbol{\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)\prime}$$);i\in\{1,\cdots,m\}\}$);
4 $\mathop{iter}\leftarrow 0$;
@@ 12       $\mathop{iter}\leftarrow\mathop{iter}+1$;
17function DistAdjust($\mathop{param}$$\boldsymbol{\Delta\widetilde{g}}_{t+1}$$\Delta\boldsymbol{g}_{t+1}^{(i)}$)
@@
[…]
@@ 25    $C_{dist}\leftarrow C_{\uparrow}$;
26
27 while $\mathop{Dist}\geq\mathop{Dist}_{max}$ && $C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}-\boldsymbol{\Delta\widetilde{g}}_{t+1}||\leq\mathop{Dist}$ do
@@ 28    $\Delta\boldsymbol{g}_{t+1}^{(i)}\leftarrow C_{dist}\cdot\Delta\boldsymbol{g}_{t+1}^{(i)}$;
The details of the two pill adjustment methods are presented as follows:
@@
Similarity-based Adjustment. As shown in Line 1--12 of Algorithm 2, we first compute the maximum cosine similarity $S_{max}$ between the normal model updates from the compromised clients and the estimated global model update in the current round. Then, we iteratively and alternately reduce the magnitude of the poison pill’s parameters with the down-scaling factor $C_{\downarrow}$, and increase the magnitude of the rest estimated global model update’s parameters with the up-scaling factor $C_{\uparrow}$, until the cosine similarity between the entire poisoned model update and the estimated global model update is greater than $S_{max}$ or the adjustment total iteration is greater than the threshold $C_{iter}$.
@@
Distance-based Adjustment. In the Distance-based Adjustment (Line 13--24 of Algorithm 2), we reuse the up-scaling factor $C_{\uparrow}$ and the down-scaling factor $C_{\downarrow}$ to adjust the magnitude of the entire poisoned model update. The intuition behind the Distance-based Adjustment is shown in Figure 3. We first calculate the maximum distance between the normal model updates from the compromised clients and the estimated global model update in the current round. We use this maximum distance $Dist_{max}$ as the threshold in the distance-based adjustment. Then, we further determine the scaling factor that should be used by applying the two scaling factors $C_{\uparrow}$ and $C_{\downarrow}$ separately to the poisoned model update $\Delta\boldsymbol{g}_{t+1}^{(i)}$. The scaling factor that reduces the distance between $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is chosen as the initial scaling factor in the subsequent iterative scaling. We stop the scaling until the distance between the $\Delta\boldsymbol{g}_{t+1}^{(i)}$ and $\boldsymbol{\Delta\widetilde{g}}_{t+1}$ is smaller than $Dist_{max}$, or the scaling factor begins to increase such distance (reach the limit of the scaling).
Appendix H Additional Experimental Configurations